OpenStudy (trwatkins1):

Which related equation can be used to solve the equation? Column A Column B 1. m + 5 = –7.3 2. m + 5 = 7.3 3. m + 7.3 = –5 4. m + 7.3 = 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5

OpenStudy (trwatkins1):

@KJSaif

KjSaif (kjsaif):

i don't get it could you screen shot it

KjSaif (kjsaif):

are u missing somthing

OpenStudy (trwatkins1):

??? idk how

KjSaif (kjsaif):

what computer do u have

OpenStudy (trwatkins1):

Match each item in Column A to an answer in Column B. 1. (1 pt) Which related equation can be used to solve the equation? Column A Column B 1. m + 5 = –7.3 2. m + 5 = 7.3 3. m + 7.3 = –5 4. m + 7.3 = 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5

OpenStudy (trwatkins1):

it comes out the same :( windowes 10 qwerty keyboard

KjSaif (kjsaif):

click on the windows menu (the window symbol)

KjSaif (kjsaif):

go to search and type snipping tool

OpenStudy (trwatkins1):

ok

KjSaif (kjsaif):

do u have it the snipping tool

OpenStudy (trwatkins1):

yep

KjSaif (kjsaif):

ok open it

OpenStudy (trwatkins1):

ok

KjSaif (kjsaif):

go to the question and and cut it

KjSaif (kjsaif):

u cut it by making a box around the question and let go of your mouse or pen

KjSaif (kjsaif):

then save as and put it in your desktop

KjSaif (kjsaif):

there u did it

OpenStudy (trwatkins1):

cool

KjSaif (kjsaif):

ok now time to answer the question

KjSaif (kjsaif):

lol

KjSaif (kjsaif):

number 1 is A

KjSaif (kjsaif):

number 2 is C

KjSaif (kjsaif):

number 3 is B

KjSaif (kjsaif):

and number 4 is B

OpenStudy (trwatkins1):

thank you btw congrats on human calculator

KjSaif (kjsaif):

thx